# An alternative form of supersymmetry with reduced cross-sections and modified experimental signatures

### 18 Dec 2023 v3

@@ I Introduction
In the present formulation, the unsolved “Gordian knot” of supersymmetry breaking (and related problems) is severed by postulating a simple primitive susy with no counterpart to (1), in which each standard model Weyl fermionic field $\psi^{r}_{f}$ is matched with a primitive (unphysical) bosonic field $\psi^{r}_{b}$;
@@
The present formulation thus trivially satisfies the Coleman-Mandula theorem Coleman-Mandula , which implies that the symmetries described by ordinary Lie algebras (with only bosonic elements) cannot mix fields of different spins. In conventional SUSY, this theorem is evaded via the extension to a graded algebra with both bosonic and fermioinc elements. Here it is satisfied because the primitive susy transformations involving the primitive bosonic fields $\psi_{b}$ manifestly do not mix different spin states.
@@
Physically acceptable bosonic fields (with excitations $a^{{\dagger}}\ket{0}$ having positive energy) are obtained by transforming the primitive $\psi_{b}$ in exactly the same way as in our previous papers, which omitted susy but led to a dark matter candidate with well-defined mass and couplings DM2021a ; DM2021b ; DM2022a ; DM2022b ; perspective . Just as before we will find that the final result is a reinterpretation of the scalar boson sectors: The original degrees of freedom in the $\psi_{b}$ are transformed into three kinds of scalar boson fields: complex Higgs-like fields $\phi$, auxiliary fields $F$, and real fields $\varphi$ of a new kind.
@@ III Experimental consequences
Corresponding to each Weyl fermion field $\psi$, there are initially two 2-component complex fields $\Phi_{s}$ (with opposite spins) having the same quantum numbers (including flavor) as $\psi$. These primitive (and unphysical) bosonic fields, with 4 total degrees of freedom, yield 4 real physical fields $\varphi_{\alpha}$ through the procedure described above. The $\varphi_{\alpha}$ (as mass eigenstates) are the final redefined sfermion fields of the present theory. They will have only the second-order gauge interactions of (64)-(66) and the second-order Higgs interactions of (48), which can now be rewritten
@@
It should be emphasized, however, that there are other processes in conventional SUSY (and shown in publications like these) which persist in the present description, with the same amplitudes as originally calculated. For example, production of squarks by direct gluon fusion and higher-order vector-boson fusion is still allowed, with the first of these depicted in Fig. 1. Calculations in our group by Tallman et al. Tallman show that the total production cross-section at the 13 TeV LHC is relatively small for a squark mass of 400 GeV– about 1 pb – and decreases with higher masses (although increasing by orders of magnitude at substantially lower masses).
@@
The search for supersymmetry is very well motivated  Baer-Tata ; Mohapatra ; Nath ; Haber-Kane ; Kane-susy ; Drees ; Baer-Barger-Tata ; Roszkowski-2018 ; Baer-Barger-2020 ; Tata-2020 ; Arkani-Hamed-Giudice ; Ellis-Olive ; 3 , and it has been a major mystery that not a single superpartner has yet been discovered. Within the present description, this may be because the cross-sections for both production and decay are reduced, and the phenomenology is very different from what has been expected (and incorporated in the analysis and simulation of events).
@@
Acknowledgement, competing interests, and data availability
The Feynman diagrams of Figs. 1-3 were prepared by Bailey Tallman. The author has no competing interests. No new data was used in this work.

### 17 Jan 2024 v4

@@ I Introduction
In the present formulation, the unsolved “Gordian knot” of supersymmetry breaking (and related problems) is severed by postulating a simple primitive susy with no counterpart to (1), in which each standard model Weyl fermionic field $\psi^{r}_{f}$ is matched with a primitive (unphysical) bosonic field $\psi^{r}_{b}$:
@@
The present formulation thus trivially satisfies the Coleman-Mandula theorem Coleman-Mandula , which implies that the symmetries described by ordinary Lie algebras (with only bosonic elements) cannot mix fields of different spins. In conventional SUSY, this theorem is evaded via the extension to a graded algebra with both bosonic and fermionic elements. Here it is satisfied because the primitive susy transformations involving the primitive bosonic fields $\psi_{b}$ manifestly do not mix different spin states.
@@
In order to obtain physically acceptable bosonic fields, we are forced to transform the primitive $\psi_{b}$ in exactly the same way as in our previous papers, which omitted susy but led to a dark matter candidate with well-defined mass and couplings DM2021a ; DM2021b ; DM2022a ; DM2022b ; perspective . Just as before we will find that the final result is a reinterpretation of the scalar boson sectors: The original degrees of freedom in the $\psi_{b}$ are transformed into three kinds of scalar boson fields: complex Higgs-like fields $\phi$, auxiliary fields $F$, and real fields $\varphi$ of a new kind.
@@ III Experimental consequences
Corresponding to each Weyl fermion field $\psi$, there are initially two complex fields $\Phi_{s}$ (with opposite spins) having the same quantum numbers (including flavor) as $\psi$. These primitive (and unphysical) bosonic fields, with 4 total degrees of freedom, are combined with their conjugate fields to yield 4 real physical fields $\varphi_{\alpha}$ through the procedure described above. The $\varphi_{\alpha}$ (as mass eigenstates) are the final redefined sfermion fields of the present theory. They will have only the second-order gauge interactions of (63)-(66) and the second-order Higgs interactions of (47), which can now be rewritten
@@
It should be emphasized, however, that there are other processes in conventional SUSY (and shown in publications like these) which persist in the present description, with the same amplitudes as originally calculated. For example, production of squarks by direct gluon fusion and higher-order vector-boson fusion is still allowed, with the first of these depicted in Fig. 1. Calculations in our group by Tallman et al. Tallman show that the total production cross-section at the 13 TeV LHC is relatively small for a squark mass of 400 GeV– about 1 pb – and decreases rapidly for higher masses (although increasing by orders of magnitude at substantially lower masses).
@@
The search for supersymmetry is very well motivated  Baer-Tata ; Mohapatra ; Nath ; Haber-Kane ; Martin ; Kane-susy ; Drees ; Baer-Barger-Tata ; Roszkowski-2018 ; Baer-Barger-2020 ; Tata-2020 ; Arkani-Hamed-Giudice ; Ellis-Olive ; 3 , and it has been a major mystery that not a single superpartner has yet been discovered. Within the present description, this may be because the cross-sections for both production and decay are reduced, and the phenomenology is very different from what has been expected (and incorporated in the analysis and simulation of events).
@@
Acknowledgement
The Feynman diagrams of Figs. 1-3 were prepared by Bailey Tallman.